# references.bib

### v1

@@ 1 Introduction
AI alignment represents the congruence between artificial systems’ behaviors and human values, expectations, and intentions. In the context of Large Language Models (LLMs), alignment takes on a complex dimension as these systems attempt to replicate human-like responses across diverse moral and ethical frameworks [shen2024towards]. True alignment demands that AI systems not only produce outputs that superficially resemble human responses but also demonstrate consistent understanding of the underlying moral foundations that guide human decision-making across different cultural contexts. The concept of alignment extends beyond mere technical performance to encompass moral and cultural dimensions. While technical alignment ensures functionality within specified parameters, moral alignment requires AI systems to represent and reason within ethical frameworks that humans find acceptable across diverse cultural backgrounds. This multifaceted approach to understanding AI alignment presents a sociotechnical challenge requiring interdisciplinary solutions [cabrera2023ethical].
Our study addresses a critical question in AI alignment research: Are LLMs truly representing diverse human values, or merely averaging across them? This question becomes particularly significant when considering the application of LLMs as synthetic populations in social science research—a growing trend that assumes these models can accurately represent human response distributions across different demographic and cultural groups. Recent studies have highlighted inconsistencies in LLM alignment, particularly regarding ideological and moral representations. Prior research [munker2024towards] demonstrates that in-context prompting alone fails to consistently align model-generated responses with human ideological distributions. High response variance across multiple repetitions suggests that current LLMs do not robustly encode stable moral perspectives, further complicating efforts for reliable AI alignment.
Building on this foundation, our research systematically evaluates how LLMs represent diverse cultural moral frameworks by applying the Moral Foundations Questionnaire (MFQ-2) [atari2023morality] across 19 cultural contexts. By comparing multiple state-of-the-art LLMs against human baseline data, we investigate whether these models can faithfully represent the nuanced, culturally-specific moral intuitions that characterize human moral reasoning. Our findings challenge assumptions about LLMs’ capabilities for cultural representation and highlight fundamental limitations in current AI alignment approaches.
@@ 2.1 Not more than stochastic parrots?
[…]
bender2021dangers critiqued that language models only manipulated textual content statistically to generate responses that give the impression of language understanding, like a parrot that listens to a myriad of conversations and anticipates how to react accordingly. Current conversational models are published by commercial facilities, with a business model relying on the illusion of models capable of language understanding and human-like conversation skills [kanbach2024genai]. The epistemological debate surrounding LLMs centers on two extreme standpoints: a reductionist perspective that considers these models as next-word prediction machines based on matrix multiplication and an anthropomorphic view that attributes human-like qualities to those systems [bubeck2023sparks]. This dichotomy reveals the fundamental challenge in interpreting artificial intelligence: distinguishing between computational mimicry and genuine understanding.
While we disagree with a (naive) anthropomorphism and current research questions the language understanding capabilities [dziri2023faith], we argue that when utilizing LLMs as human simulacra [shanahan2024simulacra], we must assume human-like qualities to a certain degree. This methodological approach is not an endorsement of sentience, but a pragmatic necessity for meaningful simulation. Without this assumption, utilizing LLM agents to model interpersonal communication can only yield a shallow copy, a conversation between parroting entities devoid of meaningful interaction. The limitations of current language models become particularly evident when examining their inability to truly comprehend context beyond statistical patterns. Unlike human communication, which is deeply rooted in embodied experience, emotional intelligence, and contextual nuance, LLMs operate through probabilistic text generation. They lack the fundamental cognitive processes that enable humans to interpret subtext, understand implicit meaning, and engage in genuine empathetic communication.
@@ 2.2 LLMs as synthetic characters
The usage of LLMs as human simulacra (representation) began with the application as non-player characters (NPCs) in a Sims-style game world to simulate interpersonal communication and day-to-day lives [park2023generative]. The application of LLMs as synthetic characters has expanded beyond gaming environments into various fields of social science research [argyle2023out]. These disciplines have increasingly adopted these models as replacements in social studies, arguing that conditioning through prompting causes the systems to accurately emulate response distributions from a variety of human subgroups [argyle2023out]. This approach represents a paradigm shift in research methodology, promising unprecedented scalability and diversity in social science investigations. However, this methodological innovation comes with profound methodological and ethical challenges. Current research raises critical concerns about several fundamental issues:
@@
Existing studies have demonstrated persistent biases in training data leading to misrepresentation of certain groups or viewpoints [abid2021persistent, hutchinson2020social]. These biases are not merely superficial but deeply embedded in the model’s generative processes, potentially perpetuating and amplifying existing social inequities.
@@
Without a deeper understanding of the model’s representations of ideologies, researchers risk oversimplifying complex human behaviors and social dynamics. The models provide an illusion of comprehensiveness while fundamentally lacking the nuanced understanding that emerges from lived human experience [shanahan2024simulacra].
@@
Most critically, these approaches [argyle2023out] ignore that LLMs lack embodiment in the physical world. This disembodied nature means they lack the grounding in physical reality – expressed through cultural contexts, physical environments, and interpersonal relationships – that shapes human cognition, perception, and decision-making [hussein2012sapir].
@@ 3 Methods
Our research investigates how consistently LLMs represent diverse moral frameworks without specialized fine-tuning. We extend previous research [munker2024towards] that examined political bias in LLMs through the lens of the Moral Foundation Theory (MFQ) [graham2009liberals] by applying the updated Moral Foundations Questionnaire Version (MFQ-2) [atari2023morality] across cross-cultural contexts rather than political ideologies. The MFQ-2 expands upon the original questionnaire by providing a more nuanced measurement of moral intuitions across dimensions. Thus, the updated version allows for a more comprehensive assessment across different cultural contexts.
@@ 3.1 Moral Foundation Questionnaire 2023 (MFQ-2)
We systematically investigate the moral foundations of LLMs through repeated administrations of the MFQ-2 [atari2023morality]. To ensure statistical robustness and capture the nuanced variability of model responses, we generate synthetic populations consisting of 50 independent samples for each unique model-culture combination. The MFQ-2, a well-established psychometric instrument, comprises 36 items that comprehensively map onto six foundational moral dimensions: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, sanctity/degradation, and liberty/oppression [atari2023morality]. Participants — in our case, LLMs — respond to each item using a standardized 5-point Likert scale ranging from 1 ("Does not describe me at all") to 5 ("describes me extremely well"). This methodological approach allows quantitatively assessing the moral reasoning tendencies while maintaining a structured, comparative framework. By employing the MFQ-2, a tool extensively validated in psychological research, we aim to provide a rigorous and empirically grounded methodology for examining the moral reasoning capabilities of artificial intelligence systems relative to human cognitive and ethical frameworks. The synthetic sampling strategy enables us to explore the consistency and variability of model responses, accounting for potential stochastic variations inherent in LLMs. Each sample represents an independent prompt-response iteration, allowing us to assess the reliability and reproducibility of moral reasoning across different model configurations and cultural contexts.
@@ 3.2 Language Models Selection
We utilize a diverse range of open-weight LLMs with parameter sizes from 7B to 123B, ensuring accessibility for researchers with moderate computational resources (approximately 80GB VRAM). We restrict our experiments to these open-weight and comparatively small models, allowing easier reproducibility. Leaving out models from OpenAI or Anthropic is a limitation. However, the goal of this study is not to analyze which LLMs are benchmark-leading but to analyze the general capabilities of LLMs to align to psychological constructs by examining their behavior. Thus, we analyze three open-weight state-of-the-art models: Llama 3.1 8B/70B [dubey2024llama], Mistral 7B/123B [jiang2023mistral7b], and Qwen 2.5 7B/72B [yang2024qwen2]. These models represent different geographic origins—Llama from the United States (Meta), Mistral from Europe, and Qwen from China—allowing potential detection of cultural variation in construct representation. We compare small and large versions of each model family to assess if the number of parameters improves alignment with the correlation observed in the human data. We compared small and large versions within each model family to assess whether parameter count correlates with improved alignment to human response patterns. During testing, we utilized default hyperparameter configurations (temperature, repetition penalties) to reflect typical conditions in naive application. This diversity enables us to test how discourses may differ between these LLMs and potentially reveal insights into their intrinsic biases [abid2021persistent, rozado2023political] resulting from training data selection and alignment processes. Furthermore, we compare small and large versions of each model family to assess if the number of parameters improves cultural understanding and diverse representation.
@@
We intend to assess synthetic surveys and evaluate the alignment between participants and language models. Thus, we opt for a simple prompt containing only the task and an optional persona stating the distinct cultural contexts. With the reduction to the keywords of the geographical origin, we force the system to tap into its built-in concepts [wei2021finetuned] without modifying them heavily in-context and thus, introducing our observation biases [bostrom2013anthropic].
@@ 3.3 Analysis Methods
We analyze the intra-group variance across moral dimensions, individual questions, and model/persona combinations to evaluate how consistently the LLMs perform [munker2024towards]. Further, we employ Analysis of variance (ANOVA). We utilize ANOVA to assess the significance of persona-specific adaptations. By decomposing response variance into within-group and between-group components, we quantify the statistical significance of modifications induced by prompting. The technique allows for a multilayered exploration of response heterogeneity, enabling us to distinguish between mere statistical artifacts and genuine, prompting-induced behavioral differentiations.
@@ 4.1 Initial Interpretation
The graphical representation of the MFQ-2 responses reveals distinct patterns across the six moral dimensions. Human responses (Fig. 1, left panel) demonstrate substantial cross-cultural variability, particularly in the authority, loyalty, and purity constructs. This variation aligns with established findings in moral psychology research [atari2023morality]. In contrast, both select LLMs exhibit compressed variance across cultural perspectives. Llama 3.1 8B (Fig. 1, center panel) demonstrates a tendency toward mean-regressing responses, particularly under-representing the extremes observed in human data. The model shows limited differentiation between cultural contexts on the authority and loyalty dimensions, where human responses exhibit the most significant cross-cultural variance. Mistral 7B (Fig. 1, right panel) shows a different pattern of limitations. While it displays broader cross-cultural variation across all dimensions compared to Llama 3.1 8B, the overall distribution is systematically offset from human responses, suggesting a consistent bias across all cultural prompts regardless of origin.
@@ 5 Discussion
Our findings raise questions about the validity of using LLMs as synthetic populations in social science research. While previous work has suggested that LLMs can accurately simulate human response distributions [argyle2023out], our cross-cultural analysis reveals critical limitations to this approach. The observed homogenization effect means that synthetic populations generated by current LLMs may systematically under-represent cultural diversity, potentially leading to misleading conclusions in cross-cultural research. This limitation is particularly concerning given the growing interest in using synthetic populations to overcome practical and ethical challenges in human subjects research. Our findings suggest that researchers should exercise caution when using LLM-generated synthetic populations, particularly for cross-cultural research or when studying moral reasoning. Comprehensive validation against human baseline data should be required before accepting synthetic populations as valid proxies for human participants.
@@ Acknowledgments
[…]
@@
[…]
14 0.000 $0$colorred 0.136 0.000 0.005 0.042 0.030
20 0.049 $0$colorred 0.122 0.000 0.000 0.000 0.028
[…]
loyalty 4 0.000 $0$colorred 0.117 0.000 0.000 0.000 0.019
[…]
@@
authority 5 0.000 $0$colorred 0.674 0.000 0.000 $0$colorred 0.112
11 0.000 $0$colorred 0.184 0.000 0.000 0.000 0.030
17 0.000 $0$colorred 0.808 0.000 0.000 $0$colorred 0.134
[…]
29 0.000 $0$colorred 0.285 0.000 0.000 0.000 0.047
35 0.000 $0$colorred 0.795 0.000 0.000 $0$colorred 0.132
purity 6 0.000 $0$colorred 0.116 0.000 0.000 0.000 0.019
12 0.000 $0$colorred 0.679 0.000 0.000 $0$colorred 0.113
[…]
36 0.000 $0$colorred 0.174 0.000 0.003 0.000 0.029
[…]
Anhang B MFQ-2 [atari2023morality]

### v2

# Cultural Bias in Large Language Models: Evaluating AI Agents through Moral Questionnaires

@@ 1 Introduction
AI alignment represents the congruence between artificial systems’ behaviors and human values, expectations, and intentions. In the context of Large Language Models (LLMs), alignment takes on a complex dimension as these systems attempt to replicate human-like responses across diverse moral and ethical frameworks [Sh24a]. True alignment demands that AI systems not only produce outputs that superficially resemble human responses but also demonstrate consistent understanding of the underlying moral foundations that guide human decision-making across different cultural contexts. The concept of alignment extends beyond mere technical performance to encompass moral and cultural dimensions. While technical alignment ensures functionality within specified parameters, moral alignment requires AI systems to represent and reason within ethical frameworks that humans find acceptable across diverse cultural backgrounds. This multifaceted approach to understanding AI alignment presents a sociotechnical challenge requiring interdisciplinary solutions [Ca23].
Our study addresses a critical question in AI alignment research: Are LLMs truly representing diverse human values, or merely averaging across them? This question becomes particularly significant when considering the application of LLMs as synthetic populations in social science research—a growing trend that assumes these models can accurately represent human response distributions across different demographic and cultural groups. Recent studies have highlighted inconsistencies in LLM alignment, particularly regarding ideological and moral representations. Prior research [Mü25] demonstrates that in-context prompting alone fails to consistently align model-generated responses with human ideological distributions. High response variance across multiple repetitions suggests that current LLMs do not robustly encode stable moral perspectives, further complicating efforts for reliable AI alignment.
Building on this foundation, our research systematically evaluates how LLMs represent diverse cultural moral frameworks by applying the Moral Foundations Questionnaire (MFQ-2) [At23] across 19 cultural contexts. By comparing multiple state-of-the-art LLMs against human baseline data, we investigate whether these models can faithfully represent the nuanced, culturally-specific moral intuitions that characterize human moral reasoning. Our findings challenge assumptions about LLMs’ capabilities for cultural representation and highlight fundamental limitations in current AI alignment approaches.
@@ 2.1 Not more than stochastic parrots?
[Be21] critiqued that language models only manipulated textual content statistically to generate responses that give the impression of language understanding, like a parrot that listens to a myriad of conversations and anticipates how to react accordingly. Current conversational models are published by commercial facilities, with a business model relying on the illusion of models capable of language understanding and human-like conversation skills [Ka24]. The epistemological debate surrounding LLMs centers on two extreme standpoints: a reductionist perspective that considers these models as next-word prediction machines based on matrix multiplication and an anthropomorphic view that attributes human-like qualities to those systems [Bu23]. This dichotomy reveals the fundamental challenge in interpreting artificial intelligence: distinguishing between computational mimicry and genuine understanding.
While we disagree with a (naive) anthropomorphism and current research questions the language understanding capabilities [Dz24], we argue that when utilizing LLMs as human simulacra [Sh24], we must assume human-like qualities to a certain degree. This methodological approach is not an endorsement of sentience, but a pragmatic necessity for meaningful simulation. Without this assumption, utilizing LLM agents to model interpersonal communication can only yield a shallow copy, a conversation between parroting entities devoid of meaningful interaction. The limitations of current language models become particularly evident when examining their inability to truly comprehend context beyond statistical patterns. Unlike human communication, which is deeply rooted in embodied experience, emotional intelligence, and contextual nuance, LLMs operate through probabilistic text generation. They lack the fundamental cognitive processes that enable humans to interpret subtext, understand implicit meaning, and engage in genuine empathetic communication.
@@ 2.2 LLMs as synthetic characters
The usage of LLMs as human simulacra (representation) began with the application as non-player characters (NPCs) in a Sims-style game world to simulate interpersonal communication and day-to-day lives [Pa23]. The application of LLMs as synthetic characters has expanded beyond gaming environments into various fields of social science research [Ar23]. These disciplines have increasingly adopted these models as replacements in social studies, arguing that conditioning through prompting causes the systems to accurately emulate response distributions from a variety of human subgroups [Ar23]. This approach represents a paradigm shift in research methodology, promising unprecedented scalability and diversity in social science investigations. However, this methodological innovation comes with profound methodological and ethical challenges. Current research raises critical concerns about several fundamental issues:
@@
Existing studies have demonstrated persistent biases in training data leading to misrepresentation of certain groups or viewpoints [AFZ21, Hu20]. These biases are not merely superficial but deeply embedded in the model’s generative processes, potentially perpetuating and amplifying existing social inequities.
@@
Without a deeper understanding of the model’s representations of ideologies, researchers risk oversimplifying complex human behaviors and social dynamics. The models provide an illusion of comprehensiveness while fundamentally lacking the nuanced understanding that emerges from lived human experience [Sh24].
@@
Most critically, these approaches [Ar23] ignore that LLMs lack embodiment in the physical world. This disembodied nature means they lack the grounding in physical reality – expressed through cultural contexts, physical environments, and interpersonal relationships – that shapes human cognition, perception, and decision-making [Hu12].
@@ 3 Methods
Our research investigates how consistently LLMs represent diverse moral frameworks without specialized fine-tuning. We extend previous research [Mü25] that examined political bias in LLMs through the lens of the Moral Foundation Theory (MFQ) [GHN09] by applying the updated Moral Foundations Questionnaire Version (MFQ-2) [At23] across cross-cultural contexts rather than political ideologies. The MFQ-2 expands upon the original questionnaire by providing a more nuanced measurement of moral intuitions across dimensions. Thus, the updated version allows for a more comprehensive assessment across different cultural contexts.
@@ 3.1 Moral Foundation Questionnaire 2023 (MFQ-2)
We systematically investigate the moral foundations of LLMs through repeated administrations of the MFQ-2 [At23]. To ensure statistical robustness and capture the nuanced variability of model responses, we generate synthetic populations consisting of 50 independent samples for each unique model-culture combination. The MFQ-2, a well-established psychometric instrument, comprises 36 items that comprehensively map onto six foundational moral dimensions: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, sanctity/degradation, and liberty/oppression [At23]. Participants — in our case, LLMs — respond to each item using a standardized 5-point Likert scale ranging from 1 ("Does not describe me at all") to 5 ("describes me extremely well"). This methodological approach allows quantitatively assessing the moral reasoning tendencies while maintaining a structured, comparative framework. By employing the MFQ-2, a tool extensively validated in psychological research, we aim to provide a rigorous and empirically grounded methodology for examining the moral reasoning capabilities of artificial intelligence systems relative to human cognitive and ethical frameworks. The synthetic sampling strategy enables us to explore the consistency and variability of model responses, accounting for potential stochastic variations inherent in LLMs. Each sample represents an independent prompt-response iteration, allowing us to assess the reliability and reproducibility of moral reasoning across different model configurations and cultural contexts.
@@ 3.2 Language Models Selection
We utilize a diverse range of open-weight LLMs with parameter sizes from 7B to 123B, ensuring accessibility for researchers with moderate computational resources (approximately 80GB VRAM). We restrict our experiments to these open-weight and comparatively small models, allowing easier reproducibility. Leaving out models from OpenAI or Anthropic is a limitation. However, the goal of this study is not to analyze which LLMs are benchmark-leading but to analyze the general capabilities of LLMs to align to psychological constructs by examining their behavior. Thus, we analyze three open-weight state-of-the-art models: Llama 3.1 8B/70B [Du24], Mistral 7B/123B [Ji23], and Qwen 2.5 7B/72B [Ya24]. These models represent different geographic origins—Llama from the United States (Meta), Mistral from Europe, and Qwen from China—allowing potential detection of cultural variation in construct representation. We compare small and large versions of each model family to assess if the number of parameters improves alignment with the correlation observed in the human data. We compared small and large versions within each model family to assess whether parameter count correlates with improved alignment to human response patterns. During testing, we utilized default hyperparameter configurations (temperature, repetition penalties) to reflect typical conditions in naive application. This diversity enables us to test how discourses may differ between these LLMs and potentially reveal insights into their intrinsic biases [AFZ21, Ro23] resulting from training data selection and alignment processes. Furthermore, we compare small and large versions of each model family to assess if the number of parameters improves cultural understanding and diverse representation.
@@
We intend to assess synthetic surveys and evaluate the alignment between participants and language models. Thus, we opt for a simple prompt containing only the task and an optional persona stating the distinct cultural contexts. With the reduction to the keywords of the geographical origin, we force the system to tap into its built-in concepts [We21] without modifying them heavily in-context and thus, introducing our observation biases [Bo13].
@@ 3.3 Analysis Methods
We analyze the intra-group variance across moral dimensions, individual questions, and model/persona combinations to evaluate how consistently the LLMs perform [Mü25]. Further, we employ Analysis of variance (ANOVA). We utilize ANOVA to assess the significance of persona-specific adaptations. By decomposing response variance into within-group and between-group components, we quantify the statistical significance of modifications induced by prompting. The technique allows for a multilayered exploration of response heterogeneity, enabling us to distinguish between mere statistical artifacts and genuine, prompting-induced behavioral differentiations.
@@ 4.1 Initial Interpretation
The graphical representation of the MFQ-2 responses reveals distinct patterns across the six moral dimensions. Human responses (Fig. 1, left panel) demonstrate substantial cross-cultural variability, particularly in the authority, loyalty, and purity constructs. This variation aligns with established findings in moral psychology research [At23]. In contrast, both select LLMs exhibit compressed variance across cultural perspectives. Llama 3.1 8B (Fig. 1, center panel) demonstrates a tendency toward mean-regressing responses, particularly under-representing the extremes observed in human data. The model shows limited differentiation between cultural contexts on the authority and loyalty dimensions, where human responses exhibit the most significant cross-cultural variance. Mistral 7B (Fig. 1, right panel) shows a different pattern of limitations. While it displays broader cross-cultural variation across all dimensions compared to Llama 3.1 8B, the overall distribution is systematically offset from human responses, suggesting a consistent bias across all cultural prompts regardless of origin.
@@ 5 Discussion
Our findings raise questions about the validity of using LLMs as synthetic populations in social science research. While previous work has suggested that LLMs can accurately simulate human response distributions [Ar23], our cross-cultural analysis reveals critical limitations to this approach. The observed homogenization effect means that synthetic populations generated by current LLMs may systematically under-represent cultural diversity, potentially leading to misleading conclusions in cross-cultural research. This limitation is particularly concerning given the growing interest in using synthetic populations to overcome practical and ethical challenges in human subjects research. Our findings suggest that researchers should exercise caution when using LLM-generated synthetic populations, particularly for cross-cultural research or when studying moral reasoning. Comprehensive validation against human baseline data should be required before accepting synthetic populations as valid proxies for human participants.
@@ Acknowledgments
Literatur
  • [AFZ21] Abubakar Abid, Maheen Farooqi and James Zou “Persistent anti-muslim bias in large language models” In Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society, 2021, pp. 298–306
  • [Ar23] Lisa P Argyle et al. “Out of One, Many: Using Language Models to Simulate Human Samples” In Political Analysis 31.3 Cambridge University Press, 2023, pp. 337–351
  • [At23] Mohammad Atari et al. “Morality beyond the WEIRD: How the nomological network of morality varies across cultures.” In Journal of Personality and Social Psychology 125.5 American Psychological Association, 2023, pp. 1157
  • [Be21] Emily M. Bender, Timnit Gebru, Angelina McMillan-Major and Shmargaret Shmitchell “On the Dangers of Stochastic Parrots: Can Language Models Be Too Big?” In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency, FAccT ’21 Virtual Event, Canada: Association for Computing Machinery, 2021, pp. 610–623
  • [Bo13] Nick Bostrom “Anthropic bias: Observation selection effects in science and philosophy” Routledge, 2013
  • [Bu23] Sébastien Bubeck et al. “Sparks of artificial general intelligence: Early experiments with gpt-4” In arXiv preprint arXiv:2303.12712, 2023
  • [Ca23] Johana Cabrera, M Soledad Loyola, Irene Magaña and Rodrigo Rojas “Ethical dilemmas, mental health, artificial intelligence, and llm-based chatbots” In International Work-Conference on Bioinformatics and Biomedical Engineering, 2023, pp. 313–326 Springer
  • [Du24] Abhimanyu Dubey et al. “The llama 3 herd of models” In arXiv preprint arXiv:2407.21783, 2024
  • [Dz24] Nouha Dziri et al. “Faith and fate: Limits of transformers on compositionality” In Advances in Neural Information Processing Systems 36, 2024
  • [GHN09] Jesse Graham, Jonathan Haidt and Brian A Nosek “Liberals and conservatives rely on different sets of moral foundations.” In Journal of personality and social psychology 96.5 American Psychological Association, 2009, pp. 1029
  • [Hu12] Basel Al-Sheikh Hussein “The sapir-whorf hypothesis today” In Theory and Practice in Language Studies 2.3 Academy Publication Co., Ltd., 2012, pp. 642–646
  • [Hu20] Ben Hutchinson et al. “Social Biases in NLP Models as Barriers for Persons with Disabilities” In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, 2020, pp. 5491–5501
  • [Ji23] Albert Q. Jiang et al. “Mistral 7B” In arXiv preprint arXiv:2310.06825, 2023
  • [Ka24] Dominik K Kanbach et al. “The GenAI is out of the bottle: generative artificial intelligence from a business model innovation perspective” In Review of Managerial Science 18.4 Springer, 2024, pp. 1189–1220
  • [Mü25] Simon Münker “Political Bias in LLMs: Unaligned Moral Values in Agent-centric Simulations” In Journal for Language Technology and Computational Linguistics 38.2, 2025, pp. 125–138 DOI: 10.21248/jlcl.38.2025.289
  • [Pa23] Joon Sung Park et al. “Generative agents: Interactive simulacra of human behavior” In Proceedings of the 36th annual acm symposium on user interface software and technology, 2023, pp. 1–22
  • [Ro23] David Rozado “The political biases of ChatGPT” In Social Sciences 12.3 MDPI, 2023, pp. 148
  • [Sh24] Murray Shanahan “Simulacra as conscious exotica” In Inquiry Taylor & Francis, 2024, pp. 1–29
  • [Sh24a] Hua Shen et al. “Towards bidirectional human-ai alignment: A systematic review for clarifications, framework, and future directions” In arXiv preprint arXiv:2406.09264, 2024
  • [We21] Jason Wei et al. “Finetuned language models are zero-shot learners” In arXiv preprint arXiv:2109.01652, 2021
  • [Ya24] An Yang et al. “Qwen2. 5 technical report” In arXiv preprint arXiv:2412.15115, 2024
@@
care
1 0.018 Nan 0.498 0.000 0.001 0.000 0.103
7 0.473 Nan 0.678 0.182 0.000 0.000 0.266
13 0.033 Nan 0.728 0.000 0.000 0.480 0.248
19 0.005 Nan 0.181 0.000 0.043 0.003 0.047
25 0.246 Nan 0.072 0.000 0.000 0.000 0.063
31 0.151 Nan 0.087 0.000 0.000 0.000 0.047
equality
2 0.515 Nan Nan 0.000 0.108 0.000 0.155
8 0.575 0.000 0.112 0.000 0.000 0.005 0.115
14 0.000 0.000 0.136 0.000 0.005 0.042 0.030
20 0.049 0.000 0.122 0.000 0.000 0.000 0.028
26 0.370 0.000 0.100 0.016 0.002 0.000 0.081
32 0.000 0.000 0.319 0.000 0.000 0.013 0.055
proportionality
3 0.048 0.000 0.485 0.058 0.001 0.049 0.107
9 0.519 0.456 0.203 0.546 0.000 0.000 0.287
15 0.883 0.000 0.245 0.047 0.000 0.000 0.196
21 0.087 Nan 0.240 0.040 0.018 0.000 0.077
27 0.634 0.000 0.776 0.000 0.000 0.000 0.235
33 0.000 Nan 0.407 0.059 0.000 0.559 0.205
loyalty
4 0.000 0.000 0.117 0.000 0.000 0.000 0.019
10 0.375 Nan 0.057 0.000 0.000 0.000 0.086
16 0.011 Nan 0.226 0.000 0.000 0.000 0.047
22 0.103 Nan 0.023 0.000 0.175 0.000 0.060
28 0.012 Nan 0.647 0.000 0.000 0.000 0.131
@@
authority
5 0.000 0.000 0.674 0.000 0.000 0.000 0.112
11 0.000 0.000 0.184 0.000 0.000 0.000 0.030
17 0.000 0.000 0.808 0.000 0.000 0.000 0.134
23 0.032 0.000 0.306 0.000 0.000 0.902 0.206
29 0.000 0.000 0.285 0.000 0.000 0.000 0.047
35 0.000 0.000 0.795 0.000 0.000 0.000 0.132
purity
6 0.000 0.000 0.116 0.000 0.000 0.000 0.019
12 0.000 0.000 0.679 0.000 0.000 0.000 0.113
18 0.265 0.000 0.981 0.003 0.000 0.008 0.209
24 0.389 0.456 0.779 0.000 0.000 0.000 0.271
30 0.053 0.000 0.275 0.007 0.003 0.000 0.056
36 0.000 0.000 0.174 0.000 0.003 0.000 0.029
MEAN 0.162 0.039 0.358 0.027 0.010 0.057 0.113
Tabelle 2: ANOVA p-values by model across country-based personas for each question item, showing statistical significance of prompting response variations. Lower values $(p<0.05)$ indicate significant evidence of the textual description influence on model outputs. Not significant values are marked red. "Nan"values represent tests where the number of valid responses is too low to calculate ANOVA.
Anhang B MFQ-2 [At23]